OpenStudy (anonymous):

integral 1 to 9 1/2x dx is log right?

OpenStudy (anonymous):

i.e 1/x dx?

OpenStudy (anonymous):

yes, integral of 1/x is log(x)

OpenStudy (anonymous):

so ie log 2x?

OpenStudy (anonymous):

No, log(x)/2 (constant stays out of integral)

OpenStudy (anonymous):

ok so 1/x is logx but since there is 2 i keep that in the denominator?

OpenStudy (anonymous):

can you help me with this

OpenStudy (anonymous):

yes, you can take it as 1/2 (constant) & stays as it is. you can take it out from integral too 1/2 integral (1/x dx)

OpenStudy (anonymous):

ok so 1/x = log x correct? and since 1/2x =logx /2?

OpenStudy (anonymous):

so i would have to log x/2 dx?

OpenStudy (anonymous):

1/2(log 9 - log 1)

OpenStudy (anonymous):

ooo ok i see...thanks

OpenStudy (anonymous):

i looked into the answer it says ln3...how would i go by caluating without calculaor?

OpenStudy (anonymous):

1/2(ln 9 - ln 1) 1/2(ln 9) (ln 1 = 0) ln sqrt(9) ln 3